# Strategic Analysis of Fair Rank-Minimizing Mechanisms with Agent Refusal Option

### v4

@@ 4 Refusal option
Let $\mathcal{A=}\left\{a_{1},a_{2},a_{3}\right\}$ and $\mathcal{O=}\left\{o_{1},o_{2},o_{3},\varnothing\right\}$. Suppose $q_{o_{1}}=q_{o_{3}}=1,$ $q_{o_{2}}=2$ and $r_{a_{1}}=(o_{1},\varnothing,o_{2},o_{3})$. Then, there are two ODSs. One is $e_{a_{1}}=(o_{1},o_{2},o_{3},\varnothing)$, which is the full extension of $a_{1}$, and the other is $d_{a_{1}}=(o_{1},o_{3},o_{2},\varnothing)$. First, suppose
$\bar{p}_{a_{2}}=(o_{1},o_{3},o_{2},\varnothing)\text{ and }\bar{p}_{a_{3}}=(o_{3},o_{1},o_{2},\varnothing).$
Then,
$f^{U}\left(r_{a_{1}},\bar{p}_{a_{2}},\bar{p}_{a_{3}}\right)=y=\left(\begin{array}[]{cccc}y_{a_{1}o_{1}}&y_{a_{1}o_{2}}&y_{a_{1}o_{3}}&y_{a_{1}\varnothing}\\ y_{a_{2}o_{1}}&y_{a_{2}o_{2}}&y_{a_{2}o_{3}}&y_{a_{2}\varnothing}\\ y_{a_{3}o_{1}}&y_{a_{3}o_{2}}&y_{a_{3}o_{3}}&y_{a_{3}\varnothing}\end{array}\right)=\left(\begin{array}[]{cccc}0&0&0&1\\ 1&0&0&0\\ 0&0&1&0\end{array}\right),$
@@
$g\left(f^{U}\left(d_{a_{1}},\bar{p}_{a_{2}},\bar{p}_{a_{3}}\right),r\right)=\left(\begin{array}[]{cccc}1/2&0&0&1/2\\ 1/2&1/2&0&0\\ 0&0&1&0\end{array}\right)$
Therefore, $a_{1}$ (strictly) prefers $g\left(f^{U}\left(d_{a_{1}},\bar{p}_{a_{2}},\bar{p}_{a_{3}}\right),r\right)$ to $g\left(f^{U}\left(r_{a_{1}},\bar{p}_{a_{2}},\bar{p}_{a_{3}}\right),r\right)$. On the other hand, there is no pair $\left(p_{a_{2}},p_{a_{3}}\right)$ of preference orders of $a_{2}$ and $a_{3}$ such that $a_{1}$ (strictly) prefers $g\left(f^{U}\left(e_{a_{1}},p_{a_{2}},p_{a_{3}}\right),r\right)$ to $g\left(f^{U}\left(r_{a_{1}},p_{a_{2}},p_{a_{3}}\right),r\right)$. This is because, by Proposition 3 and $q_{o_{1}}+q_{o_{2}}=3=\left|\mathcal{A}\right|$, for any $\left(p_{a_{2}},p_{a_{3}}\right)\in\mathcal{P}^{2}$ and any $\bar{o}\in\mathcal{O}$,
@@
We explain why $d_{a_{1}}$ strategically dominates $e_{a_{1}}$. In this example, $q_{o_{1}}+q_{o_{3}}=2<3=q_{o_{1}}+q_{o_{2}}$. That is, the copies of $o_{1}$ and $o_{2}$ are sufficient to allocate all agents, but those of $o_{1}$ and $o_{3}$ are insufficient. Thus, revealing $o_{2}$ as a second-best type does not increase the probability that $a_{1}$ is assigned to $o_{1}$. On the other hand, revealing $o_{3}$ as a second-best type increases it, when $o_{3}$ is the first-best type of another agent. Thus, revealing the full extension may be strategically dominated by another ODS when an unacceptable type (in this case $o_{3}$) is less preferable than the other (in this case $o_{2}$) but the number of copies of the former is less than that of the latter.
Generally, we have the following result.
Lemma 4
Suppose that there is $a\in\mathcal{A}$ whose true preference satisfies
$R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)=R\left(o^{\prime},r_{a}\right)-1<R\left(o^{\prime\prime},r_{a}\right),\text{ }$
$\sum\limits_{\bar{o}:R\left(\bar{o},r_{a}\right)<R\left(\varnothing,r_{a}\right)}q_{\bar{o}}+q_{o^{\prime}}\geq\left|\mathcal{A}\right|,\text{ and}$
$\sum\limits_{\bar{o}:R\left(\bar{o},r_{a}\right)<R\left(\varnothing,r_{a}\right)}q_{\bar{o}}+q_{o^{\prime\prime}}<\left|\mathcal{A}\right|,$
[…]
$R\left(\varnothing,r_{a}\right)=R\left(o^{\prime},e_{a}\right)=R\left(o^{\prime\prime},d_{a}\right),$
with the refusal option.
Note that Lemma 4 does not contradict a result of Featherstone (2020, Proposition 13). In his model, he assumes that the agents have no information on the capacities of object types. Revealing the full extension can be strategically dominated by another ODS only when the capacities of the unacceptable types are different and agents know them. Thus, the strategic domination in Lemma 4 does not occur in the low-information situation discussed by Featherstone (2020).
Next, we show that if agents adopt an ODS, then the result may be wasteful and thus inefficient.
@@
By Theorem 2, for an agent $a$ with $r_{a}$ that satisfy (3) for some $o,o^{\prime}\in\mathcal{O}$, an ODS dominates truth-telling. Moreover, by Proposition 1, in the same case, the strategy may make the assignment wasteful and thus inefficient. In an ODS, an agent reveals actually unacceptable types as acceptable types. Then, the agent may be assigned to an actually unacceptable type and refuse it. Since the type may be desirable for another agent, in that case, the assignment is wasteful.
By Theorems 1 and 2 and Proposition 3, if the uniform rank-minimizing mechanism is used, the agents may have a strong incentive to manipulate their preferences, and these manipulations may result in an inefficient assignment. Thus, next, we consider another fair rank-minimizing mechanism such that, contrary to $f^{U}$, if it is used as the mechanism in the second stage, then for any agent, truth-telling is not strategically dominated by any other strategy.
@@
for agent $a,$ $R\left(o,p_{a}\right)=1$ and $R\left(\varnothing,p_{a}\right)\geq 3,$
@@
and if $p$ does not satisfy all of them, then
@@
To understand this mechanism, we reconsider Example 1. First, suppose $q_{o_{1}}=q_{o_{2}}=1$. If $\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)=\left(\psi^{\prime},\psi,\psi\right)$, $\left(\psi,\psi^{\prime},\psi\right)$ or $\left(\psi,\psi,\psi^{\prime}\right)$, then this satisfies (i), (ii) and (iii). In these cases, a student who chooses $\psi^{\prime}$ is corresponding to $a,$ another student is corresponding to $a^{\prime}$, and $o_{1}$ is corresponding to $o$. The rank-minimizing mechanism $f$ introduced in Example 1 is the modified uniform rank-minimizing mechanism, because a student who chooses $\psi^{\prime}$ obtains $o_{1}$ with probability 00 and another student obtains $o_{1}$ with probability $1$
On the other hand, suppose $q_{o_{1}}=1$ and $q_{o_{2}}\geq 2$. Even if $\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)=\left(\psi^{\prime},\psi,\psi\right)$, $\left(\psi,\psi^{\prime},\psi\right)$ or $\left(\psi,\psi,\psi^{\prime}\right)$, this does not satisfy (ii), because in this case, $p_{a^{\prime}}=\psi$ and $l=3.$ Then,
@@
Therefore, (ii) is not satisfied with $p_{a^{\prime}}=\psi$.
@@
The strategic issues, mentioned in Theorem 2, do not arise under this mechanism $f^{M}$.
First, we explain why the strategic issue mentioned in Example 2 does not arise in the modified uniform rank-minimizing mechanism. If $r_{a}=p_{a^{\prime}}$ where $p_{a^{\prime}}$ is defined in (ii), then $\left(d_{a},p_{-a}\right)$ satisfies (i), (ii) and (iii) where $d_{a}\in\mathcal{D}_{a}$. This implies that for some $p_{-a}$, if agent $a$ adopts an ODS instead of truth-telling and $f^{M}$ is used, then $a$ is worse off.
To see this fact, we revisit Example 2. In this example, $\left(d_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right)$ satisfies (i), (ii) and (iii), but $\left(r_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right)$ does not. Thus,
$f^{M}\left(r_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right)$ $=$ $f^{U}\left(r_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right)=\left(\begin{array}[]{ccc}1/3&0&2/3\\ 1/3&0&2/3\\ 1/3&0&2/3\end{array}\right),$
$f^{M}\left(d_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right)$ $=$ $\left(\begin{array}[]{ccc}0&1&0\\ 1/2&0&1/2\\ 1/2&0&1/2\end{array}\right),$
$g\left(f^{M}\left(d_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right)\right)$ $=$ $\left(\begin{array}[]{ccc}0&0&1\\ 1/2&0&1/2\\ 1/2&0&1/2\end{array}\right).$
This implies that $a_{1}$ may be worse off by revealing $d_{a_{1}}$ instead of $r_{a_{1}}$, even if $a_{1}$ refuses any unacceptable objects. Therefore, $r_{a_{1}}$ is not strategically dominated by $d_{a_{1}}$, if $f^{M}$ is used.
@@
Suppose that the modified uniform rank-minimizing mechanism is used. For any $r_{a}\in\mathcal{P}$, there is no $p_{a}\in\mathcal{P}$ that strategically dominates $r_{a}$ with (or without) the refusal option.
By adopting $f^{M}$, any ODS can fail, because increasing the number of acceptable objects may lead the agent to be unassigned acceptable objects in their true preference. Thus, in some profiles of other agents’ (revealed) preferences, truth-telling is better than adopting any ODS. However, $f^{M}$ may lead to other strategic problems. That is, if $f^{M}$ is adopted instead of $f^{U}$, an agent may have an incentive to reduce the number of acceptable object types.
@@
Let $R\left(\varnothing,r_{a}\right)=k$. Suppose $k\geq 3$; that is, $\varnothing$ is neither the first-best type nor the second-best type for $r_{a}$. Thus, we let $r_{a}^{1}=o$ and $r_{a}^{2}=o^{\prime}$. For $p_{-a}$, there are $q_{o}$ agents represented by $a^{\prime}$ whose revealed preference order is such that $p_{a^{\prime}}^{k^{\prime}}=r_{a^{\prime}}^{k^{\prime}}$ for all $k^{\prime}=1,\cdots,k-2$ and $p_{a^{\prime}}^{k-1}=\varnothing$. We assume
$q_{r_{a}^{1}}+\cdots+q_{r_{a}^{k-2}}<\left|\mathcal{A}\right|\text{.}$
Moreover, for $p_{-a}$ and any other agent represented by $a^{\prime\prime}$, $R\left(\varnothing,r_{a^{\prime\prime}}\right)=1$. In this case, since $\left(r_{a},p_{-a}\right)$ satisfies (i), (ii) and (iii),
$\left(f^{M}\left(r_{a},p_{-a}\right)\right)_{ao}$ $=$ 0, ($0,\text{ }\left(f^{M}\left(r_{a},p_{-a}\right)\right)_{ao^{\prime}}=1,$
$\left(f^{U}\left(r_{a},p_{-a}\right)\right)_{ao}$ $=$ $\frac{q_{o}}{q_{o}+1},$
$\left(f^{U}\left(r_{a},p_{-a}\right)\right)_{ao^{\prime}}$ $=$ $\frac{1}{q_{o}+1}.$
On the other hand, we let $p_{a}=p_{a^{\prime}}$. Since $\left(p_{a},p_{-a}\right)$ does not satisfy all of (i), (ii) and (iii),
$\left(f^{M}\left(p_{a},p_{-a}\right)\right)_{ao}$ $=$ $\left(f^{U}\left(p_{a},p_{-a}\right)\right)_{ao}=\frac{q_{o}}{q_{o}+1},$
$\left(f^{M}\left(p_{a},p_{-a}\right)\right)_{ao^{\prime}}$ $=$ $\left(f^{U}\left(p_{a},p_{-a}\right)\right)_{ao^{\prime}}=\frac{1}{q_{o}+1}.$
Thus, in this case, if $f^{M}$ is used, then $a$ has an incentive to choose $p_{a}$. In other words, in this case, $a$ has an incentive to reveal by reducing its truly acceptable types. On the other hand, if $f^{U}$ is used, then $a$ has no such an incentive.
Of course, this does not mean that truth-telling is strategically dominated revealing some reduction of $r_{a}$. However, agents may have such incentives even if they do not have the refusal option.

### v5

@@ 4 Refusal option
Let $\mathcal{A=}\left\{a_{1},a_{2},a_{3}\right\}$ and $\mathcal{O=}\left\{o_{1},o_{2},o_{3},\varnothing\right\}$. Let $\gamma=(o_{1},o_{2},o_{3},\varnothing),$ $\gamma^{\prime}=(o_{1},o_{3},o_{2},\varnothing)$ and $\gamma^{\prime\prime}=(o_{3},o_{1},o_{2},\varnothing).$ Suppose $q_{o_{1}}=q_{o_{3}}=1,$ $q_{o_{2}}=2$ and $r_{a_{1}}=(o_{1},\varnothing,o_{2},o_{3})$. Then, there are two ODSs. One is $e_{a_{1}}=\gamma$, which is the full extension of $a_{1}$, and the other is $d_{a_{1}}=\gamma^{\prime}$. In this example, we assume $\left(p_{a_{2}},p_{a_{3}}\right)=\left(r_{a_{2}},r_{a_{3}}\right);$ that is, $a_{2}$ and $a_{3}$ always adopt truth-telling.
First, suppose $p_{a_{2}}=\gamma^{\prime}$ and $p_{a_{3}}=\gamma^{\prime\prime}.$ Then,
$f^{U}\left(r_{a_{1}},\gamma^{\prime},\gamma^{\prime\prime}\right)=y=\left(\begin{array}[]{cccc}y_{a_{1}o_{1}}&y_{a_{1}o_{2}}&y_{a_{1}o_{3}}&y_{a_{1}\varnothing}\\ y_{a_{2}o_{1}}&y_{a_{2}o_{2}}&y_{a_{2}o_{3}}&y_{a_{2}\varnothing}\\ y_{a_{3}o_{1}}&y_{a_{3}o_{2}}&y_{a_{3}o_{3}}&y_{a_{3}\varnothing}\end{array}\right)=\left(\begin{array}[]{cccc}0&0&0&1\\ 1&0&0&0\\ 0&0&1&0\end{array}\right),$
@@
$g\left(f^{U}\left(d_{a_{1}},\gamma^{\prime},\gamma^{\prime\prime}\right),r\right)=\left(\begin{array}[]{cccc}1/2&0&0&1/2\\ 1/2&1/2&0&0\\ 0&0&1&0\end{array}\right)\text{.}$
Therefore, $a_{1}$ (strictly) prefers $g\left(f^{U}\left(d_{a_{1}},\gamma^{\prime},\gamma^{\prime\prime}\right),r\right)$ to $g\left(f^{U}\left(r_{a_{1}},\gamma^{\prime},\gamma^{\prime\prime}\right),r\right)$. On the other hand, there is no pair $\left(p_{a_{2}},p_{a_{3}}\right)$ of preference orders of $a_{2}$ and $a_{3}$ such that $a_{1}$ (strictly) prefers $g\left(f^{U}\left(e_{a_{1}},p_{a_{2}},p_{a_{3}}\right),r\right)$ to $g\left(f^{U}\left(r_{a_{1}},p_{a_{2}},p_{a_{3}}\right),r\right)$. This is because, by $q_{o_{1}}+q_{o_{2}}=3=\left|\mathcal{A}\right|$, for any $\left(p_{a_{2}},p_{a_{3}}\right)\in\mathcal{P}^{2}$ and any $\bar{o}\in\mathcal{O}$,
@@
We explain why $d_{a_{1}}$ strategically dominates $e_{a_{1}}$. In this example, $q_{o_{1}}+q_{o_{3}}=2<3=q_{o_{1}}+q_{o_{2}}$. That is, the copies of $o_{1}$ and $o_{2}$ are sufficient to allocate all agents, but those of $o_{1}$ and $o_{3}$ are insufficient. Thus, revealing $o_{2}$ as a second-best type does not increase the probability that $a_{1}$ is assigned to $o_{1}$. On the other hand, revealing $o_{3}$ as a second-best type increases it, when $o_{3}$ is the first-best type of another agent. This is because, to minimize the total rank, $o_{3}$ must be assigned to an agent who ranks it as their first-best type. However, since $a_{1}$ ranks $o_{2}$ and $\varnothing$ lower, minimizing the total rank requires increasing the probability that $o_{1}$ is allocated to $a_{1}$. Thus, revealing the full extension may be strategically dominated by another ODS when an unacceptable type (in this case $o_{3}$) is less preferable than the other (in this case $o_{2}$) but the number of copies of the former is less than that of the latter.
Note that this fact does not contradict a result of Featherstone (2020, Proposition 13). In their model, they assume that the agents have no information on the capacities of object types. Revealing the full extension can be strategically dominated by another ODS only when the capacities of the unacceptable types are different and agents know them. Thus, the strategic domination in Example 3 does not occur in the low-information situation discussed by Featherstone (2020).
Next, we show that if some agent adopts an ODS, then the result may be wasteful and thus inefficient.
@@
By Theorem 2, for an agent $a$ with $r_{a}$ that satisfy (3) for some $o,o^{\prime}\in\mathcal{O}$, an ODS dominates truth-telling. Moreover, by Proposition 3, in the same case, the strategy may make the assignment wasteful. In an ODS, an agent reveals actually unacceptable types as acceptable types. Then, the agent may be assigned to an actually unacceptable type and refuse it. Since the type may be desirable for another agent, in that case, the assignment is wasteful.
By Theorems 1 and 2 and Proposition 3, if the uniform rank-minimizing mechanism is employed, the agents may have a strong incentive to manipulate their preferences, and these manipulations may result in an inefficient assignment. Thus, next, we consider another fair rank-minimizing mechanism such that, contrary to $f^{U}$, if it is employed as the mechanism in the second stage, then for any agent, truth-telling is not strategically dominated by any other strategy.
@@
there is one agent represented by $a$ such that $R\left(o,p_{a}\right)=1$ and $R\left(\varnothing,p_{a}\right)\geq 3,$
@@
and otherwise; that is, if $p$ does not satisfy all of them, then
@@
To understand this mechanism, we reconsider Example 1. First, suppose $q_{o_{1}}=q_{o_{2}}=1$. If $\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)=\left(\alpha^{\prime},\alpha,\alpha\right)$, $\left(\alpha,\alpha^{\prime},\alpha\right)$ or $\left(\alpha,\alpha,\alpha^{\prime}\right)$, then $\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)$ satisfies (i), (ii) and (iii). In these cases, the student who chooses $\alpha^{\prime}$ is corresponding to $a,$ another student is corresponding to $a^{\prime}$, and $o_{1}$ is corresponding to $o$. The rank-minimizing mechanism $f^{\ast}$ introduced in Example 1 is the modified uniform rank-minimizing mechanism, because a student who chooses $\alpha^{\prime}$ obtains $o_{1}$ with probability 00 and another student obtains $o_{1}$ with probability $1$
On the other hand, suppose $q_{o_{1}}=1$ and $q_{o_{2}}\geq 2$. Even if $\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)=\left(\alpha^{\prime},\alpha,\alpha\right)$, $\left(\alpha,\alpha^{\prime},\alpha\right)$ or $\left(\alpha,\alpha,\alpha^{\prime}\right)$, this does not satisfy (ii). This is because in this case, $p_{a^{\prime}}=\alpha$ and $l=3,$ but
@@
Therefore, (ii) is not satisfied with $p_{a^{\prime}}=\alpha$ in the case of $q_{o_{1}}=1$ and $q_{o_{2}}\geq 2$.
@@
The strategic issues, mentioned in Theorems 1 and 2, do not arise under this mechanism $f^{M}$.
First, we explain why the strategic issue mentioned in Example 2 does not arise in the modified uniform rank-minimizing mechanism. If $r_{a}=p_{a^{\prime}}$ where $p_{a^{\prime}}$ is defined in (ii), then $\left(d_{a},p_{-a}\right)$ satisfies (i), (ii) and (iii) where $d_{a}\in\mathcal{D}_{a}$. This implies that for some $p_{-a}$, if agent $a$ adopts an ODS instead of truth-telling and $f^{M}$ is employed, then $a$ is worse off.
To see this fact, we revisit Example 2. In this example, $\left(d_{a_{1}},p_{a_{2}},p_{a_{3}}\right)=\left(\beta,\beta^{\prime\prime},\beta^{\prime\prime}\right)$ satisfies (i), (ii) and (iii), but $\left(r_{a_{1}},p_{a_{2}},p_{a_{3}}\right)=\left(\beta^{\prime\prime},\beta^{\prime\prime},\beta^{\prime\prime}\right)$ does not. Thus,
$f^{M}\left(r_{a_{1}},\beta^{\prime\prime},\beta^{\prime\prime}\right)$ $=$ $f^{U}\left(r_{a_{1}},\beta^{\prime\prime},\beta^{\prime\prime}\right)=\left(\begin{array}[]{ccc}1/3&0&2/3\\ 1/3&0&2/3\\ 1/3&0&2/3\end{array}\right),$
$f^{M}\left(d_{a_{1}},\beta^{\prime\prime},\beta^{\prime\prime}\right)$ $=$ $\left(\begin{array}[]{ccc}0&1&0\\ 1/2&0&1/2\\ 1/2&0&1/2\end{array}\right),$
$g\left(f^{M}\left(d_{a_{1}},\beta^{\prime\prime},\beta^{\prime\prime}\right)\right)$ $=$ $\left(\begin{array}[]{ccc}0&0&1\\ 1/2&0&1/2\\ 1/2&0&1/2\end{array}\right).$
This implies that $a_{1}$ may be worse off by revealing $d_{a_{1}}$ instead of $r_{a_{1}}$, even if $a_{1}$ refuses any unacceptable objects. Therefore, $r_{a_{1}}$ is not strategically dominated by $d_{a_{1}}$, if $f^{M}$ is employed.
@@
Suppose that the modified uniform rank-minimizing mechanism is employed. For any $r_{a}\in\mathcal{P}$, there is no $p_{a}\in\mathcal{P}$ that strategically dominates $r_{a}$ with (or without) the refusal option.
By adopting $f^{M}$, any ODS can fail because increasing the number of acceptable objects may lead the agent to be unassigned to truly acceptable objects, depending on the strategy profile of the other agents. Thus, in some profiles of other agents’ (revealed) preferences, truth-telling is better than adopting any ODS. However, $f^{M}$ may lead to other strategic problems. That is, if $f^{M}$ is adopted instead of $f^{U}$, an agent may have an incentive to reduce the number of acceptable object types.
@@
Let $\mathcal{A=}\left\{a_{1},a_{2}\right\}$ and $\mathcal{O=}\left\{o_{1},o_{2},\varnothing\right\}$. Let $\delta=\left(o_{1},o_{2},\varnothing\right)$ and $\delta^{\prime}=\left(o_{1},\varnothing,o_{2}\right)$. We assume $q_{o}=1$. Suppose $r_{a_{1}}=\delta$ and $p_{a_{2}}=\delta$. In this case, since $\left(\delta(=r_{a_{1}}),\delta\right)$ does not satisfy all of (i), (ii) and (iii),
$\left(f^{M}\left(\delta,\delta\right)\right)_{a_{1}o_{1}}$ $=$ $\left(f^{M}\left(\delta,\delta\right)\right)_{a_{1}o_{2}}=\frac{1}{2},$
$\left(f^{U}\left(\delta,\delta\right)\right)_{a_{1}o_{1}}$ $=$ $\left(f^{U}\left(\delta,\delta\right)\right)_{a_{1}o_{2}}=\frac{1}{2}.$
On the other hand, since $\left(\delta^{\prime}(\neq r_{a_{1}}),\delta\right)$ satisfies all of (i), (ii) and (iii),
$\left(f^{M}\left(\delta^{\prime},\delta\right)\right)_{a_{1}o_{1}}$ $=$ 1, ($1,\text{ }\left(f^{U}\left(\delta^{\prime},\delta\right)\right)_{a_{1}o_{2}}=0,$
$\left(f^{U}\left(\delta^{\prime},\delta\right)\right)_{a_{1}o_{1}}$ $=$ $\left(f^{U}\left(\delta^{\prime},\delta\right)\right)_{a_{1}o_{2}}=\frac{1}{2}.$
Thus, in this case, if $f^{M}$ is employed, then $a_{1}$ has an incentive to choose $\delta^{\prime}$ instead of $r_{a_{1}}=\delta$. In other words, in this case, $a_{1}$ has an incentive to reveal by reducing its truly acceptable types. On the other hand, if $f^{U}$ is employed, then $a_{1}$ has no such an incentive.
Of course, this does not mean that truth-telling is strategically dominated revealing some reduction of $r_{a_{1}}$. However, agents may have such incentives even if they do not have the refusal option.